# Burnside type results for fusion categories

### v10

@@ 6.3. On the support $J_{{(L,\;\mathcal{S})}}$
There is a dual version of Theorem 6.5 and its Corollary 6.6:
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then $x_{i}\in{G{(H,\mathcal{B})}}$ if and only $i\in\mathcal{I}_{{{\widehat{(H,\mathcal{B})}}}_{ad}}$.
Dually one obtains also the following:
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then
(6.3) $\lambda_{{}_{{{\widehat{(H,\mathcal{B})}}}_{ad}}}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}{\widehat{E}}_{i}.$
Proof of Theorem 1.7 and Theorem 1.8.
@@
If ${(H,\mathcal{B})}$ is Burnside then by Corollary 4.5,
$\big{(}\prod_{j=1}^{m}\mu_{j}\big{)}^{2}=\sum_{x_{i}\in G(H,B)}{\widehat{E}}_{i}.$
Corollary 6.8 implies Equation (1.4). Conversely, if Equation (1.4) holds then
$\prod_{j=1}^{m}\mu_{j}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}\pm{\widehat{E}}_{i}$
Evaluating both sides at any ${\frac{x_{i}}{d_{i}}}$, one obtains the Burnside property. ∎
By duality one obtains directly Theorem 1.8.
6.4. On the sub-hypergroup generated by $P$
Proposition 6.9.
For any two sub-hypergroups ${(L,\;\mathcal{S})},{(K,\mathcal{T})}$ of an ARN-hypergroup ${(H,\mathcal{B})}$, then $\mathcal{J}_{{}_{{(L,\;\mathcal{S})}\vee{(K,\mathcal{T})}}}={\mathcal{J}_{{}_{{(K,\mathcal{T})}}}}\cap{\mathcal{J}_{{}_{{(L,\;\mathcal{S})}}}}$.
@@ 6.4. On the sub-hypergroup generated by $P$
We prove that $\lambda_{{}_{{(L,\;\mathcal{S})}\vee{(K,\mathcal{T})}}}={\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}$. Then Equation (6.1) implies the intersection result. Note that for all $s\in\mathcal{S}$ and $t\in\mathcal{T}$, $st{\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}=d_{s}d_{t}{\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}$. Lemma 5.3 shows that for any constituent $u$ of $st$ then $u{\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}=d_{u}{\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}$ also. Thus the same relation holds for any standard element of ${(L,\;\mathcal{S})}\vee{(K,\mathcal{T})}$. ∎
Corollary 6.10.
For any two sub-hypergroups ${(L,\;\mathcal{S})},{(K,\mathcal{T})}$ of an ARN-hypergroup ${(H,\mathcal{B})}$, then ${(L,\;\mathcal{S})}\subseteq{(K,\mathcal{T})}$ if and only if ${\mathcal{J}_{{}_{{(K,\mathcal{T})}}}}\subseteq{\mathcal{J}_{{}_{{(L,\;\mathcal{S})}}}}$.
Proof.
If ${(L,\;\mathcal{S})}\subseteq{(K,\mathcal{T})}$ then ${\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}={\lambda_{{}_{{(K,\mathcal{T})}}}}$ and therefore equation (6.1) implies ${\mathcal{J}_{{}_{{(K,\mathcal{T})}}}}\subseteq{\mathcal{J}_{{}_{{(L,\;\mathcal{S})}}}}$. The converse follows from Proposition 6.9. ∎
Corollary 6.11.
For any two elements $P,Q\in{(H,\mathcal{B})}_{+}$ of an ARN-hypergroup ${(H,\mathcal{B})}$, then $<P>\subseteq<Q>$ if and only if ${\ker_{{}_{{(H,\mathcal{B})}}}}(Q)\subseteq{\ker_{{}_{{(H,\mathcal{B})}}}}(P)$.
Recall $P:=\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}$ for any ${(H,\mathcal{B},\mu_{1})}$.
Proposition 6.12.
In any abelian RN-hypergroup ${(H,\mathcal{B})}$ one has
$<P^{2}>={{(H,\mathcal{B})}_{ad}}.$
Proof.
One can prove first that ${\ker_{{}_{{(H,\mathcal{B})}}}}(P^{2})={\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$. Indeed $\big{|}\psi(P)\big{|}=\prod_{i=1}^{m}\big{|}\psi({\frac{x_{i}}{d_{i}}})\big{|}\leq 1$. Thus, if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(P)$, then $\big{|}\psi({\frac{x_{i}}{d_{i}}})\big{|}=1$, i.e $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$, by Proposition 6.1. Conversely, if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$ then $\psi({\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}})=\big{|}\psi({\frac{x_{i}}{d_{i}}})\big{|}^{2}=1$. Thus, if $i\neq i^{*}$ then grouping together $x_{i}$ and $x_{i^{*}}$ in $P$, one obtains $\psi({\frac{x_{i}}{d_{i}}})\psi(\frac{x_{{i^{*}}}}{{d_{i^{*}}}})=\big{|}\psi({\frac{x_{i}}{d_{i}}})^{2}\big{|}=1$. Thus,
$\psi(P)=\prod_{i={i^{*}}}\psi({\frac{x_{i}}{d_{i}}}).$
Note that if $i=i^{*}$ then $\psi({\frac{x_{i}}{d_{i}}})=\pm 1$, since in this case $\psi(x_{i})=\psi(x_{i^{*}})\in\mathbb{R}$. Thus, $\psi(P)=\pm 1$ and $\psi(P^{2})=1$ for all $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$.
It follows that $\lambda_{<P^{2}>}=\lambda_{<I(1)>}$ by Theorem 5.7, and so $<P^{2}>={{(H,\mathcal{B})}_{ad}}$. ∎
Corollary 6.13.
In any ARN-hypergroup ${(H,\mathcal{B})}$ one has
${{(H,\mathcal{B})}_{ad}}\subseteq<P>.$
Corollary 6.14.
Let ${(H,\mathcal{B})}$ be an abelian RN hypergroup. Then the following holds:
  1. (1)
    ${{(H,\mathcal{B})}_{ad}}=\mathbb{C}$ if and only if ${\widehat{(H,\mathcal{B})}}$ is pointed.
  2. (2)
    ${{(H,\mathcal{B})}_{ad}}={(H,\mathcal{B})}$ if and only if ${\widehat{(H,\mathcal{B})}}$ is perfect.
@@
A simple non-pointed abelian RN-hypergroup ${(H,\mathcal{B})}$ has a perfect dual.
@@
By simplicity ${{(H,\mathcal{B})}_{ad}}={(H,\mathcal{B})}$, and therefore ${\widehat{(H,\mathcal{B})}}$ is perfect. ∎
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup with ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ RN. Then the following holds:
@@
    ${\widehat{(H,\mathcal{B})}}_{ad}={\widehat{(H,\mathcal{B})}}$ if and only if ${(H,\mathcal{B})}$ is perfect.

### v11

@@ 6.3. On the support ${\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}$
Here are dual versions of Theorem 6.5 and Corollary 6.6:
Theorem 6.7.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then $x_{i}\in{G{(H,\mathcal{B})}}$ if and only $i\in\mathcal{I}_{{{\widehat{(H,\mathcal{B})}}}_{ad}}$.
Corollary 6.8.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then
(6.3) $\lambda_{{}_{{{\widehat{(H,\mathcal{B})}}}_{ad}}}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}\widehat{F}_{i}.$
6.4. Proof of Theorems 1.7 and 1.10
Recall Definition 6.3, which defines the support $\mathcal{I}_{{(L,\mathcal{S})}}\subset\mathcal{I}$ of a sub-hypergroup ${(L,\mathcal{S})}$ within an abelian hypergroup ${(H,\mathcal{B})}$. For the sake of brevity, we define $|L|:=n(L,\mathcal{S},\mu_{1})$.
Theorem 6.9.
Let ${(L,\mathcal{S})}$ be a sub-hypergroup of an abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$. With the above notations, we have:
(6.4) $\sum\limits_{x_{i}\in\mathcal{S}}\widehat{F}_{i}=\frac{|L|}{|H|}\big{(}\sum\limits_{j\in\mathcal{I}_{{(L,\mathcal{S})}}}{\widehat{h}}_{j}\mu_{j}\big{)}.$
Proof.
Applying $\mathcal{F}$ to Equation (6.1) we have that
[…]
On the other hand, from Equation (2.13) we have ${\lambda_{{}_{{(L,\mathcal{S})}}}}=\frac{1}{|L|}\big{(}\sum\limits_{x_{i}\in\mathcal{S}}h_{i^{*}}d_{i^{*}}x_{i}\big{)}$ and therefore:
[…]
Comparing (6.5) and (6.6), we obtain:
$\sum\limits_{x_{i}\in\mathcal{S}}\widehat{F}_{i}=|L|\big{(}\sum\limits_{j\in\mathcal{I}_{(L,\mathcal{S})}}\frac{\mu_{j}}{n_{j}}\big{)}.$
According to Equation (2.11), we have ${\widehat{h}}_{j}=\frac{|H|}{n_{j}}$, thus concluding the proof. ∎
Proof of Theorem 1.7: By applying Corollary 4.5, we get that
$\big{(}\prod_{j\in\mathcal{I}}\mu_{j}\big{)}^{2}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}\widehat{F}_{i}.$
On the other hand, by utilizing Theorem 6.9 with $\mathcal{S}=G{(H,\mathcal{B})}$, we arrive at the desired conclusion.
Proof of Theorem 1.10: To derive the result through duality from Theorem 1.7, we need to consider the normalized basis $(\frac{x_{i}}{d_{i}})_{i\in\mathcal{I}}$. Recall that by definition, $\langle x_{i}x_{i^{*}},x_{1}\rangle^{-1}=h_{i}$. Therefore, $\langle\frac{x_{i}}{d_{i}}\frac{x_{i^{*}}}{d_{i^{*}}},x_{1}\rangle^{-1}=d_{i}d_{i^{*}}h_{i}$. Consequently, $d_{i}d_{i^{*}}h_{i}\frac{x_{i}}{d_{i}}=d_{i^{*}}h_{i}x_{i}$, which precisely matches the summand on the RHS.
6.5. Proof of Theorems 1.8 and 1.11.
Proof.
By Corollary 4.5, ${(H,\mathcal{B})}$ is Burnside if and only if
$\big{(}\prod_{j\in\mathcal{I}}\mu_{j}\big{)}^{2}=\sum_{x_{i}\in G(H,B)}\widehat{F}_{i},$
if and only if Equation (1.4) holds, by Corollary 6.8, (the dual version of) Equation (2.12) and Lemma 2.14.∎
By duality, we obtain directly Theorem 1.11.
6.6. On the sub-hypergroup generated by $P$
Lemma 6.10.
For any sub-hypergroup ${(L,\mathcal{S})}$ of an ARN-hypergroup ${(H,\mathcal{B})}$, then ${\mathcal{I}}_{{(L,\mathcal{S})}}=\bigcap_{s\in\mathcal{S}}{\mathcal{I}}_{s}$. Thus, $s\in\mathcal{S}$ if and only if $s\lambda_{{(L,\mathcal{S})}}=d_{s}\lambda_{{(L,\mathcal{S})}}$.
Proof.
Note that ${(L,\mathcal{S})}=\langle x\rangle$, with $x=\sum_{s\in\mathcal{S}}s\in{(H,\mathcal{B})}_{+}$. Recall that ${\mathcal{I}}_{x}$ is the index set of ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)$. By Theorem 5.7, ${\mathcal{I}_{x}}={\mathcal{I}_{\langle x\rangle}}$, and by Lemma 5.5, ${\mathcal{I}_{x}}=\bigcap_{s\in\mathcal{S}}{\mathcal{I}_{s}}$. The first sentence follows. Recall that $x=\sum_{j\in\mathcal{I}}\mu_{j}(x)F_{j}$, thus $x\lambda_{{(L,\mathcal{S})}}=\sum_{j\in\mathcal{I}_{{(L,\mathcal{S})}}}\mu_{j}(x)F_{j}$. But the first sentence means that $j\in{\mathcal{I}}_{{(L,\mathcal{S})}}$ if and only if $\mu_{j}(s)=\mathrm{FPdim}(s)$, for all $s\in\mathcal{S}$. So if $s\in\mathcal{S}$ then $s\lambda_{{(L,\mathcal{S})}}=d_{s}\lambda_{{(L,\mathcal{S})}}$. Regarding the converse, we deduce by positivity in Equation (5.2) that $sS\subseteq S$. Specifically, since $s=sx_{1}\in sS$, it follows that $s\in S$. ∎
In particular, for an ARN-hypergroup with $\mu_{1}=\mathrm{FPdim}$, then ${\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}=\{j\in{\mathcal{I}}\text{ such that }\mu_{j}|_{(L,\mathcal{S})}=\mathrm{FPdim}\}$. Let ${(L,\mathcal{S})},{(K,\mathcal{T})}$ be two sub-hypergroups of an ARN-hypergroup ${(H,\mathcal{B})}$. Let ${(L,\mathcal{S})}\vee{(K,\mathcal{T})}$ be the sub-hypergroup generated by ${(L,\mathcal{S})}$ and ${(K,\mathcal{T})}$.
@@ 6.6. On the sub-hypergroup generated by $P$
With above notations, $\mathcal{I}_{{}_{{(L,\mathcal{S})}\vee{(K,\mathcal{T})}}}={\mathcal{I}_{{}_{{(K,\mathcal{T})}}}}\cap{\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}$. In other words, $\lambda_{{(L,\mathcal{S})}\vee{(K,\mathcal{T})}}=\lambda_{{(L,\mathcal{S})}}\lambda_{{(K,\mathcal{T})}}$.
Proof.
Let ${\mathcal{S}}\vee{\mathcal{T}}$ denotes the basis of ${(L,\mathcal{S})}\vee{(K,\mathcal{T})}$. By Lemmas 6.10 and 5.3,
${\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}\cap{\mathcal{I}_{{}_{{(K,\mathcal{T})}}}}=\bigcap_{a\in{\mathcal{S}}\cup{\mathcal{T}}}{\mathcal{I}}_{a}$ =$=\{j\in{\mathcal{I}}\ |\ \mu_{j}\in\ker_{{(H,\mathcal{B})}}(a),\ \forall a\in{\mathcal{S}}\cup{\mathcal{T}}\}$
=$=\{j\in{\mathcal{I}}\ |\ {\mathcal{S}}\cup{\mathcal{T}}\subset\ker_{{\widehat{(H,\mathcal{B})}}}(\mu_{j})\}$
=$=\{j\in{\mathcal{I}}\ |\ {\mathcal{S}}\vee{\mathcal{T}}\subset\ker_{{\widehat{(H,\mathcal{B})}}}(\mu_{j})\}$
=$=\{j\in{\mathcal{I}}\ |\ \mu_{j}\in\ker_{{(H,\mathcal{B})}}(a),\ \forall a\in{\mathcal{S}}\vee{\mathcal{T}}\}$
$=\bigcap_{a\in{\mathcal{S}}\vee{\mathcal{T}}}{\mathcal{I}}_{a}=\mathcal{I}_{{}_{{(L,\mathcal{S})}\vee{(K,\mathcal{T})}}}.\qed$
Proposition 6.12.
For any two sub-hypergroups ${(L,\mathcal{S})},{(K,\mathcal{T})}$ of an ARN-hypergroup ${(H,\mathcal{B})}$, then ${(L,\mathcal{S})}\subseteq{(K,\mathcal{T})}$ if and only if ${\mathcal{I}_{{}_{{(K,\mathcal{T})}}}}\subseteq{\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}$.
Proof.
By Lemma 6.10, if ${(L,\mathcal{S})}\subseteq{(K,\mathcal{T})}$ then ${\mathcal{I}_{{}_{{(K,\mathcal{T})}}}}\subseteq{\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}$. By Proposition 6.11, if ${\mathcal{I}_{{}_{{(K,\mathcal{T})}}}}\subseteq{\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}$, then ${\lambda_{{}_{{(L,\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}={\lambda_{{}_{{(K,\mathcal{T})}}}}$. Pick $s\in\mathcal{S}$, then
$s{\lambda_{{}_{{(K,\mathcal{T})}}}}=s{\lambda_{{}_{{(L,\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}=d_{s}{\lambda_{{}_{{(L,\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}=d_{s}{\lambda_{{}_{{(K,\mathcal{T})}}}},$
thus $s\in\mathcal{T}$, by Lemma 6.10. ∎
@@
For any two elements $P,Q\in{(H,\mathcal{B})}_{+}$ of an ARN-hypergroup ${(H,\mathcal{B})}$, then $\langle P\rangle\subseteq\langle Q\rangle$ if and only if ${\ker_{{}_{{(H,\mathcal{B})}}}}(Q)\subseteq{\ker_{{}_{{(H,\mathcal{B})}}}}(P)$.
For any ${(H,\mathcal{B},\mu_{1})}$, recall that $P:=\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}}$.
Proposition 6.14.
For any abelian RN-hypergroup ${(H,\mathcal{B})}$ then
$\langle P^{2}\rangle={{(H,\mathcal{B})}_{ad}}.$
@@
By Definition 6.2, ${{(H,\mathcal{B})}_{ad}}:=\langle I(1)\rangle$, so we are reduced to show that $\mathcal{I}_{\langle P^{2}\rangle}=\mathcal{I}_{\langle I(1)\rangle}$, so (by Theorem 5.7) that $\mathcal{I}_{P^{2}}=\mathcal{I}_{I(1)}$ i.e. ${\ker_{{}_{{(H,\mathcal{B})}}}}(P^{2})={\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$.
Note that $\big{|}\psi(P)\big{|}=\prod_{i\in\mathcal{I}}\big{|}\psi({\frac{x_{i}}{d_{i}}})\big{|}\leq 1$. Thus, if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(P)$, then $\big{|}\psi({\frac{x_{i}}{d_{i}}})\big{|}=1$, i.e $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$, by Proposition 6.1. Thus, ${\ker_{{}_{{(H,\mathcal{B})}}}}(P)\subseteq{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$, and the same argument shows that ${\ker_{{}_{{(H,\mathcal{B})}}}}(P^{2})\subseteq{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$. Conversely, if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$ then $\psi({\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}})=\big{|}\psi({\frac{x_{i}}{d_{i}}})\big{|}^{2}=1$. Thus, if $i\neq i^{*}$ then, grouping together $x_{i}$ and $x_{i^{*}}$ in $P$, we obtain $\psi({\frac{x_{i}}{d_{i}}})\psi(\frac{x_{{i^{*}}}}{{d_{i^{*}}}})=\big{|}\psi({\frac{x_{i}}{d_{i}}})^{2}\big{|}=1$. Thus,
$\psi(P)=\prod_{i={i^{*}}}\psi({\frac{x_{i}}{d_{i}}}).$
Note that if $i=i^{*}$, then $\psi({\frac{x_{i}}{d_{i}}})=\pm 1$, since in this case $\psi(x_{i})=\psi(x_{i^{*}})\in\mathbb{R}$. Thus, if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$, then $\psi(P)=\pm 1$ and $\psi(P^{2})=1$, i.e. $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(P^{2})$. ∎
Corollary 6.15.
For any ARN-hypergroup ${(H,\mathcal{B})}$, then ${{(H,\mathcal{B})}_{ad}}\subseteq\langle P\rangle$.
Corollary 6.16.
Let ${(H,\mathcal{B})}$ be an ARN hypergroup. The following holds:
  1. (1)
    ${{(H,\mathcal{B})}_{ad}}=\mathbb{C}$ if and only if ${(H,\mathcal{B})}$ is pointed,
  2. (2)
    ${{(H,\mathcal{B})}_{ad}}=\mathbb{C}$ if and only if $^{\widehat{(H,\mathcal{B})}}$ is pointed,
  3. (3)
    ${{(H,\mathcal{B})}_{ad}}={(H,\mathcal{B})}$ if and only if $^{\widehat{(H,\mathcal{B})}}$ is perfect.
@@
First, recall that ${{(H,\mathcal{B})}_{ad}}=\langle I(1)\rangle$, so ${{(H,\mathcal{B})}_{ad}}=\mathbb{C}$ if and only if $I(1)=n{(H,\mathcal{B})}x_{1}$, if and only if ${(H,\mathcal{B})}$ is pointed, as $I(1)=\sum_{i\in\mathcal{I}}h_{i}x_{i}x_{i^{*}}$. Next, by Equation (5.2), ${{(H,\mathcal{B})}_{ad}}=\mathbb{C}$ if and only if $\lambda_{{{(H,\mathcal{B})}_{ad}}}=\lambda_{\mathbb{C}}=x_{1}=\operatorname{id}=\sum_{i\in\mathcal{I}}F_{i},$ if and only if $G{\widehat{(H,\mathcal{B})}}=\widehat{\mathcal{B}}$, by Equation 6.2, meaning that $^{\widehat{(H,\mathcal{B})}}$ is pointed. Similary, ${{(H,\mathcal{B})}_{ad}}={(H,\mathcal{B})}$ if and only if $\lambda_{{{(H,\mathcal{B})}_{ad}}}=F_{1}$, if and only if $G{\widehat{(H,\mathcal{B})}}=\{\mu_{1}\}$, meaning that $^{\widehat{(H,\mathcal{B})}}$ is perfect. ∎
@@
A simple non-pointed ARN-hypergroup ${(H,\mathcal{B})}$ has a perfect dual.
@@
By simplicity ${{(H,\mathcal{B})}_{ad}}=\mathbb{C}$ or ${(H,\mathcal{B})}$. By non-pointed assumption and Corollary 6.16, ${{(H,\mathcal{B})}_{ad}}\neq\mathbb{C}$, therefore ${{(H,\mathcal{B})}_{ad}}={(H,\mathcal{B})}$, and so $^{\widehat{(H,\mathcal{B})}}$ is perfect by Corollary 6.16. ∎
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup with $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ RN. Then the following holds:
  1. (1)
    ${\widehat{(H,\mathcal{B})}}_{ad}=\mathbb{C}$ if and only if $^{\widehat{(H,\mathcal{B})}}$ is pointed,
@@
    $^{\widehat{(H,\mathcal{B})}}_{ad}={\widehat{(H,\mathcal{B})}}$ if and only if ${(H,\mathcal{B})}$ is perfect.